# Assessing the viability of non-light water reactor concepts for electricity and heat generation in decarbonized energy systems

### v1

@@ Appendix D Energy vectors and technologies in the system model
The right column of Fig. 4 lists the final energy demands the model considers. Section LABEL:demFlex details how the model represents the flexibility of final demand. The final electricity demand corresponds to residential, service, and industry appliances. Next, there is a final demand for space and process heating. The model further splits process heat into three temperature levels: low temperature up to 100°C, medium temperature from 100°C to 500°C, and high temperature above 500°C. To cover the demand for space heat, the model can invest in solar thermal heating, boilers and heat-pumps fuelled by electricity, the connection to a district heating network, and boilers fuelled by biomass or synthetic fuels. The same options are available for low temperature heat, except for solar thermal heating; medium and high temperatures preclude heat-pumps and district heating. Even if a technology can provide a particular heat type, its capacity can still be constrained. The potential for district heating and ground-source heat-pumps in space heating reflects settlement structures. District heating can only supply heat in cities, towns, and suburbs; ground-source heat-pumps in rural areas and half the demand in towns and suburbs. In process heating, today’s shares of district heating and electric boilers are an upper limit since it is indeterminate whether their temperature level is sufficient to supply a larger share of demand. Technologies for space heat and low-temperature heat utilize heat storage, as section LABEL:demFlex elaborates.
@@
Appendix E Spatial resolution and infrastructure in the system model
@@
We use a transport instead of flow formulation to model the grid operation since previous research found this simplification to be sufficiently accurate [163, cited by 115]. In line with the same source, transmission losses amount to 5% and 3% per 1,000 km for HVAC and HVDC grids, respectively. The distance used is the distance between the geographic centers of each zone.
Figure 13: Spatial resolution and grid connections for hydrogen [gogoke_flexible_2023]

### v2

@@ Appendix E Energy Vectors and Technologies in the System Model
The right column of Fig. 2 lists the final energy demands the model considers. The final electricity demand corresponds to residential, service, and industry appliances. Next, there is a final demand for space and process heating. The model further splits process heat into three temperature levels: low temperature up to 100°C, medium temperature from 100°C to 500°C, and high temperature above 500°C. To cover the demand for space heat, the model can invest in solar thermal heating, boilers and heat-pumps fuelled by electricity, the connection to a district heating network, and boilers fuelled by biomass or synthetic fuels. The same options are available for low temperature heat, except for solar thermal heating; medium and high temperatures preclude heat-pumps and district heating. Even if a technology can provide a particular heat type, its capacity can still be constrained. The potential for district heating and ground-source heat-pumps in space heating reflects settlement structures. District heating can only supply heat in cities, towns, and suburbs; ground-source heat-pumps in rural areas and half the demand in towns and suburbs. In process heating, today’s shares of district heating and electric boilers are an upper limit since it is indeterminate whether their temperature level is sufficient to supply a larger share of demand. Technologies for space heat and low-temperature heat utilize heat storage.
@@
Appendix F Spatial Resolution and Infrastructure in the System Model
@@
We use a transport instead of flow formulation to model the grid operation since previous research found this simplification to be sufficiently accurate [190, cited by 87]. In line with the same source, transmission losses amount to 5% and 3% per 1,000 km for HVAC and HVDC grids, respectively. The distance used is the distance between the geographic centers of each zone.
Figure 15: Spatial resolution and grid connections for hydrogen [9]